OpenStudy (mitu12):

simplify (3a^2)(5a^6)

OpenStudy (mathstudent55):

Do you know how to multiply to powers with the same base? For example, what is \(4^6 \times 4^{10} = \)

OpenStudy (mathstudent55):

I meant "two powers"

OpenStudy (mitu12):

16^16

OpenStudy (mitu12):

yes i think the answer will be 15a8

OpenStudy (mitu12):

to the 8th power

OpenStudy (mathstudent55):

Yes, for your problem, answer is \(15a^8\); you add the exponents. For my example above, you are close, but just add the exponents and leave the base the same it was. \(4^6 \times 4^{10} = 4^{16} \) See, the 4 stays as it is, it does not become 16.

OpenStudy (mitu12):

thank you i understand it now

OpenStudy (mathstudent55):

Great. You're welcome.